# Evaluating the Fitness of Ontologies for the Task of Question Generation

### v1

@@ Abstract
Ontology-based question generation is an important application of semantic-aware systems that enables the creation of large question banks for diverse learning environments. The effectiveness of these systems, both in terms of the calibre and cognitive difficulty of the resulting questions, depends heavily on the quality and modelling approach of the underlying ontologies, making it crucial to assess their fitness for this task. To date, there has been no comprehensive investigation into the specific ontology aspects or characteristics that affect the question generation process. Therefore, this paper proposes a set of requirements and task-specific metrics for evaluating the fitness of ontologies for question generation tasks in pedagogical settings. Using the ROMEO methodology, a structured framework for deriving task-specific metrics, an expert-based approach is employed to assess the performance of various ontologies in Automatic Question Generation (AQG) tasks, which is then evaluated over a set of ontologies. Our results demonstrate that ontology characteristics significantly impact the effectiveness of question generation, with different ontologies exhibiting varying performance levels. This highlights the importance of assessing ontology quality with respect to AQG tasks.
Keywords:
Ontology Evaluation Question Generation Ontologies
@@ 1 Introduction
Research on the use of semantic models within Automatic Question Generation (AQG) has greatly advanced the field of educational assessment. Ontologies, with their structured representation of knowledge, introduce an additional layer of semantic understanding which allows AQG systems to generate questions modelled at a semantic level [7], thereby surpassing syntax- and vocabulary-focused approaches that operate at a shallower level [17]. Furthermore, semantic-aware AQG models that exploit ontological elements such as classes and properties, have demonstrated superior generalisability across different domains and question formats, compared to machine learning-based approaches [3]. Factors such as the size of the ontology, its hierarchical organisation, domain coverage, and defined semantic relations play a key role in determining the properties of the output [23, 37, 44]; e.g., the linguistic issues in the questions generated have been found to be linked to how concepts are modelled within the ontology [23]. Yet there has been no systematic investigation into the type of ontology features (or how they are quantified) that can enhance or hinder pedagogical question generation.
@@
The choice of ontology is crucial when generating and evaluating questions using AQG systems. Although it is common practice within the AQG research community to develop experimental ontologies [2, 5, 13, 24, 40, 42, 43], this has been criticised for potentially affecting the generalisability and validity of the frameworks and their question patterns, due to them being purpose-built and not representative of real-world applications [5]. Reusing existing ontologies would address these concerns by ensuring the frameworks had been tested on widely accepted and standardised knowledge structures, thereby enhancing their applicability and reliability across diverse contexts. However, identifying and selecting suitable ontologies that are fit for purpose is complex, as this requires the expertise, intuition, and domain knowledge of an ontology specialist. Similarly, when creating new ontologies, frameworks are needed that can guide construction and be used to evaluate the final product. Approaches such as Application or Task-based evaluation can assess the suitability and fitness of an ontology for a specific task, by evaluating the extent to which an ontology conforms to the task requirements [9, 18, 19, 26, 33]. However, for this to be applied to AQG, the requirements (and evaluation metrics) for question generation must first be specified, before they can be used to evaluate how well an ontology supports them.
@@
This paper focuses on identifying the requirements and ontology evaluation metrics specifically designed to assess the suitability of an ontology for AQG. By using the ROMEO methodology [46] (a requirements-oriented framework designed for evaluating ontologies in task-specific contexts), a core set of requirements and associated metrics have been identified and evaluated over a set of ontologies. This expert-led approach combines theoretical and practical perspectives, by integrating domain ontologies into a question-generation model and deriving relevant metrics based on observed performance. Section 2 provides an overview of AQG together with the challenges involved in evaluating ontologies, and the methodology adopted to determine a set of targetted evaluation requirements for AQG is presented in Section 3. The resulting requirements, evaluation metrics and evaluations are detailed in Section 4, before concluding in Section 5.
@@ 2 Background and Related Work
Ontology-based question generation, whereby a domain ontology is used as the knowledge source to generate questions, typically involves the modelling of generic question types using a set of templates containing specific Resource Description Framework (RDF) patterns. These are used to generate different types of questions using ontological elements such as concepts [6, 13], various types of properties [37] and individuals [13]. Several of the most common question formats are illustrated in Table 1, together with the corresponding RDF pattern requirements for each question type (which vary both in terms of the number of triple patterns they require and the types of elements they include). Papasalouros & Chatzigiannakou [29] classify ontology-based questions based on the type of ontological components targetted by the generation strategies: class-based, property-based and terminology-based (Table 1). Class membership questions assess the learner’s understanding of class-instance relationships, such as identifying the class to which a specific entity belongs [1, 5, 13], whereas Property-based questions focus on testing the learner’s knowledge of an entity’s properties and their values, such as the attributes or relationships of a given entity [1, 30, 36]. Terminology-based questions explore hierarchical relationships of an entity within the ontology, such as subclass and superclass associations [5, 11, 29]. In addition to these question types, Annotation-based questions [5, 36] leverage annotations in the ontology to generate questions about descriptions or the definitions of terms. These generation strategies result in questions that have a one-to-one mapping between the number of relevant triples and questions generated, and as such each pattern template only tests the learners’ knowledge about a single fact.
@@
Question Generation Strategy RDF Pattern Approaches
@@
Beyond these foundational types, other common question types incorporate additional triples or target special types of semantic relationships, such as Multiple Choice Questions (MCQs) and multi-entity questions. MCQs require not only the generation of a question but also the creation of a set of distractors that are plausible yet distinct from the correct answer [2, 5, 13, 25, 37, 41]. Multi-entity questions synthesise knowledge from multiple interconnected entities, requiring the learner to have a deeper understanding of the relationships and dependencies between various ontological elements [25, 35, 37]. These advanced question types are particularly valuable for evaluating higher-order thinking and comprehension [12, 25, 37], and whilst MCQs and multi-entity questions may overlap with the previous categories in their use of classes, properties, or instances, their defining characteristic is the integration of multiple axioms to form complex questions and the construction of distractors, that introduces an additional cognitive challenge.
Domain ontologies, which represent the knowledge of a specific subject area, form the primary input for question generation models. Many studies rely on hand-crafted ontologies developed specifically for evaluation purposes [2, 5, 13, 24, 40, 42, 43], which frequently focus on the knowledge of a particular course [5, 24], or represent broader knowledge of a general domain [2, 13, 24, 40, 42, 43]. Consequently, such ontologies tend to be small in scope and shallow in structure, capturing only the most general aspects of the domain. This limited depth and scale may constrain the capabilities of question generation models and hinder a comprehensive assessment of their effectiveness. Additionally, the process of building domain-specific ontologies from scratch is time-intensive and requires specialised expertise, posing a challenge to the acceleration of research in this area. Furthermore, whether a new ontology is engineered or an existing one is reused, it is still important to quantitatively assess if it is fit for the task of question generation through the use of appropriate ontology evaluation metrics.
@@
As ontologies are used across diverse application contexts, it is crucial that robust evaluation methods are used to assess them before their adoption within new systems. Ontology evaluation involves a comprehensive examination of their structure, alignment with domain requirements, and use for specific tasks [18, 19]. Such methods can be broadly categorised into four main categories [9, 19, 45]:
Gold Standard evaluation:
where the ontology is compared against a gold standard, which is an established and authoritative ontology or knowledge base covering the same domain. This evaluates the ontology’s coverage by identifying gaps, inconsistencies, or redundancies.
User-centred evaluation:
this involves obtaining and analysing feedback from end-users or domain experts who evaluate the ontology.
Data-driven evaluation:
directly compares an ontology to existing data, such as textual documents related to the domain it models.
Application or task-based evaluation:
this is performed when an ontology is used for a specific application or task. This evaluation assesses the effectiveness of the ontology in the context of an application or task.
When reusing an ontology for a specific task, it is imperative to evaluate its suitability and alignment with the task’s requirements. A requirement is defined as “an expression of desired behaviour[32], with ontology requirements specifically denoting the desired quality or competency of an ontology within the context of an application [46]. Given the diversity of ontology requirements across applications, a thorough analysis of the target application needs is essential before ontology reuse. Once these requirements are identified, they can be operationalised qualitatively, through expert judgments or user feedback, or quantitatively, employing measurable metrics (i.e. measures). Although various evaluation measures for ontologies have been proposed, selecting appropriate measures should be guided by the application’s requirements rather than relying on an arbitrary set of metrics [16, 27, 31, 46].
@@ 3 Requirement-Oriented Methodology for AQG
In order to determine an appropriate set of ontology-evaluation criteria for AQG, it is necessary to first identify those ontology characteristics that are both relevant to the generation of good quality questions, as well as being measurable. ROMEO (Requirements-Oriented Methodology for Evaluating Ontologies) is a requirement-oriented approach for identifying a set of unique requirements and characteristics for a task domain (in this context, AQG), that can then be aligned to a set of relevant evaluation metrics [46]. It was originally adapted from the Goal-Question-Metrics (GQM) framework [10], designed to derive quality measures by linking them to the goals of a given software application. By determining the requirements of the AQG process through a user-based study, suitable metrics can be identified that quantify the degree to which an ontology can effectively fulfil its intended purpose; i.e. facilitating the generation of high calibre and cognitively challenging questions.
An expert-led study was conducted to determine a number of ontology requirements given a set of generated questions, based on the following stages:
  1. 1.
    A small number of existing domain ontologies were selected for the purpose of generating questions (Table 2); namely, an astronomy-based ontology focused on the solar system,111 https://anonymous.4open.science/r/SolarSystemOntology-F64E/ an ontology representing U.S. geographical data, [^2] http://www.cs.utexas.edu/users/ml/geo.html and an ontology capturing intuitive knowledge about African wildlife [20].
  2. 2.
    Several question generation patterns were utilised to generate questions by instantiating templates with ontology elements (Table 1). This resulted in the generation of 3482 questions [^3] For MCQs, only unique question stems were considered; repeated questions with varying answer choices were excluded to avoid redundancy. across all three ontologies (Table 4).
  3. 3.
    An expert familiar with AQG was asked to assess the quality of questions, by completing a questionnaire containing a mix of open and closed questions. Two criteria were used to assess the questions; effectiveness and appropriateness (see below). Furthermore, they were asked to identify any ontology-related requirements that may impact the quality of the generation process.
  4. 4.
The ontologies listed in Table 2 were chosen as they exhibit different characteristics and thus could be used to examine how their structural differences influence the question generation outcome; such as an emphasis on taxonomic modelling or hierarchical classifications and conceptual relationships. Two of these were also populated, thus providing rich factual and instance-based representations. The questions that were presented to the expert were generated using an AQG approach [4] that utilised the RDF patterns in Table 1 within a fixed textual template. Additional post-processing techniques, such as verbalisation (which are typically carried out to enhance the generation results), were intentionally excluded to maintain consistency in the structure and format of the generated questions across different ontologies, and to ensure that the expert-led comparison of question quality and effectiveness was not influenced by such techniques. [^4] Example questions appearing later in this paper have been paraphrased for clarity.
Table 2: Basic statistics of the ontologies utilised in study.
@@
An expert reviewed the question specifications detailed in Table 1 and evaluated the questions generated from each of the three ontologies; with the aim of identifying the specific ontological features that contributed to the variations in the questions. The two evaluation criteria employed by Leo et al. [25] were used when assessing each ontology and its corresponding questions:
@@
  • The question quality (appropriateness) is assessed based on how pedagogically appropriate the questions are in order to assess learners at different knowledge levels, by ensuring that there is a range of questions from basic to more advanced understanding of the domain.
As the evaluation task involved an extensive qualitative analysis of a large pool of questions generated from three diverse ontologies, a single expert was recruited by this study to assess the feasibility of the proposed approach, as well as to understand critical aspects such as the time and effort needed by experts for such evaluation.
4 Response Analysis and Requirement Derivation
This section presents the outcomes of the expert-led study, by describing the observations made by the expert, the requirements derived from these observations, and proposed metrics that measure the requirements. The proposed evaluation metrics build upon well-established measures [16, 28, 38, 39] that have been adapted to align with the specific requirements of the AQG task. The model used for defining metrics formally specifies the ontology schema and the knowledge base structures. An ontology schema represents the class and property definitions (i.e. T-box) and is denoted by the 5-tuple $\mathcal{O}=\{\mathcal{C},\mathcal{P},\mathcal{H}^{\mathcal{C}},prop,\mathcal{A}\}$, whereas a knowledge base (i.e. A-box) represents the instantiated ontological elements (individuals) and is denoted by the 5-tuple $\mathcal{KB}=\{\mathcal{O},\mathcal{I},L,inst^{\mathcal{C}},inst^{\mathcal{P}}\}$. The meaning of the symbols used in these definitions are given in Table 3. Whilst this model is similar to others proposed for various ontology scenarios [28], it has been adapted to be more pertinent to the metrics that focus on AQG tasks.
Table 3: Symbols used by the metrics following the nomenclature used in [38, 39].
@@ 4 Response Analysis and Requirement Derivation
$\mathcal{H}^{\mathcal{C}}$ a concept hierarchy, which is a directed, transitive relation, and relates to the taxonomy. $\mathcal{H}^{\mathcal{C}}(\mathcal{C}_{1},\mathcal{C}_{2})$ states that $\mathcal{C}_{1}$ is a sub-concept of $\mathcal{C}_{2}$ (i.e. $\mathcal{C}_{1}\sqsubseteq\mathcal{C}_{2}$).
$prop$: $\mathcal{P}\to\mathcal{C}\times\mathcal{C}$: A function mapping properties that link pairs of concepts. $prop(\mathcal{P})=(\mathcal{C}_{1},\mathcal{C}_{2})$ can also be expressed as $\mathcal{P}(\mathcal{C}_{1},\mathcal{C}_{2})$
$\mathcal{A}$ a set containing annotation properties of the type rdfs:comment
@@
$inst^{\mathcal{C}}$: $\mathcal{C}\to 2^{\mathcal{I}}$ (concept instantiation) maps concepts to their instances. $inst^{\mathcal{C}}(\mathcal{C})=\mathcal{I}$ may also be written as $\mathcal{C}(\mathcal{I})$.
$inst^{\mathcal{P}}$: $\mathcal{P}\to 2^{\mathcal{I}\times Ob}$ maps properties to their instantiated objects (i.e. $Ob=\mathcal{I}\cup L$). $inst^{\mathcal{P}}(\mathcal{I},Ob)=\mathcal{I}$ may also be written as $\mathcal{P}(\mathcal{I},Ob)$.

### v2

@@ Abstract
Ontology-based question generation is an important application of semantic-aware systems that enables the creation of large question banks for diverse learning environments. The effectiveness of these systems, both in terms of the calibre and cognitive difficulty of the resulting questions, depends heavily on the quality and modelling approach of the underlying ontologies, making it crucial to assess their fitness for this task. To date, there has been no comprehensive investigation into the specific ontology aspects or characteristics that affect the question generation process. Therefore, this paper proposes a set of requirements and task-specific metrics for evaluating the fitness of ontologies for question generation tasks in pedagogical settings. Using the ROMEO methodology (a structured framework used for identifying task-specific metrics), a set of evaluation metrics have been derived from an expert assessment of questions generated by a question generation model. To validate the proposed metrics, we apply them to a set of ontologies previously used in question generation to illustrate how the metric scores align with and complement findings reported in earlier studies. The analysis confirms that ontology characteristics significantly impact the effectiveness of question generation, with different ontologies exhibiting varying performance levels. This highlights the importance of assessing ontology quality with respect to Automatic Question Generation (AQG) tasks.
@@ 1 Introduction
Research on the use of semantic models within Automatic Question Generation (AQG) has greatly advanced the field of educational assessment. Ontologies, with their structured representation of knowledge, introduce an additional layer of semantic understanding which allows AQG systems to generate questions modelled at a semantic level [7], thereby surpassing syntax- and vocabulary-focused approaches that operate at a shallower level [17]. Furthermore, semantic-aware AQG models that exploit ontological elements (such as classes and properties) have demonstrated superior generalisability across different domains and question formats, compared to machine learning-based approaches [3]. Factors such as the size of the ontology, its hierarchical organisation, domain coverage, and defined semantic relations also play a key role in determining the properties of the output [23, 30, 37, 44]; e.g., the linguistic issues in the questions generated have been found to be linked to how concepts are modelled within the ontology [23]. Yet, to date, there has been no systematic investigation into the type of ontology features that can enhance or hinder pedagogical question generation.
@@
The choice of ontology is crucial when generating and evaluating questions using AQG systems. Although it is common practice within the AQG research community to develop experimental ontologies [2, 5, 13, 24, 43, 42, 40], this has been criticised for potentially affecting the generalisability and validity of the frameworks and their question patterns, due to them being purpose-built and not representative of real-world applications [4, 5]. Reusing existing ontologies would address these concerns by ensuring that the frameworks had been tested on widely accepted and standardised knowledge structures, thereby enhancing their applicability and reliability across diverse contexts. However, identifying and selecting suitable ontologies that are fit for purpose is complex, as this requires the expertise, intuition, and domain knowledge of an ontology specialist. Similarly, when creating new ontologies, frameworks are needed that can guide construction and be used to evaluate the final product. Approaches such as Application or Task-based evaluation can assess the suitability and fitness of an ontology for a specific task, by evaluating the extent to which an ontology conforms to the task requirements [9, 18, 19, 26, 33]. However, for this to be applied to AQG, the requirements (and evaluation metrics) for question generation must first be specified, before they can be used to evaluate how well an ontology supports them.
@@
This paper focuses on identifying the requirements and ontology evaluation metrics that are specifically designed to assess the suitability of an ontology for AQG. By using the ROMEO methodology [45] (a requirements-oriented framework designed for evaluating ontologies in task-specific contexts), a core set of requirements and associated metrics have been identified and evaluated over a set of ontologies. This expert-led approach combines theoretical and practical perspectives, by integrating domain ontologies into a question-generation model and deriving the relevant metrics based on their observed performance. Section 2 provides an overview of AQG together with the challenges involved in evaluating ontologies, and the methodology adopted to determine a set of targeted evaluation requirements for AQG is presented in Section 3. The resulting requirements, evaluation metrics and evaluations are detailed in Section 4, before validating the metrics across a separate set of ontologies used by other AQG studies, to correlate the scores with known question properties in Section 5. The paper concludes in Section 6.
@@ 2 Background and Related Work
Ontology-based question generation, whereby a domain ontology is used as the knowledge source to generate questions, typically involves the modelling of generic question types using a set of templates containing specific Resource Description Framework (RDF) patterns. This generates different types of questions using ontological elements such as concepts [6, 13], various types of properties [37] and individuals [13]. Several of the most common question formats are illustrated in Table 1, with the corresponding RDF pattern requirements that vary both in terms of the number of triple patterns they require and the types of elements they include. [^1] Input ontologies and the resulting questions for each generation strategy are available at https://github.com/SamahAlkhuzaey/Onto_eval_AQG.git
Papasalouros & Chatzigiannakou [29] classify ontology-based questions based on the type of ontological components targeted by the generation strategies: Class-based, Property-based and Terminology-based (Table 1). Class-based questions assess the learner’s understanding of class-instance (i.e. class membership) relationships, such as identifying the class to which a specific entity belongs [1, 5, 13], whereas Property-based questions focus on testing the learner’s knowledge of an entity’s properties and their values, such as the attributes or relationships of a given entity [1, 30, 36]. Terminology-based questions explore hierarchical relationships of an entity within the ontology, such as subclass and superclass associations [5, 11, 29]. Finally, Annotation-based questions [5, 36] leverage annotations in the ontology to generate questions about descriptions or the definitions of terms. These generation strategies result in questions that have a one-to-one mapping between the number of relevant triples and questions generated, and as such, each pattern template only tests the learners’ knowledge about a single fact.
Beyond these foundational types, other common question types incorporate additional triples or target special types of semantic relationships, such as Multiple Choice Questions (MCQs) and multi-entity questions. MCQs require not only the generation of a question but also the creation of a set of distractors that are plausible yet distinct from the correct answer [2, 5, 13, 25, 37, 41]. Multi-entity questions synthesise knowledge from multiple interconnected entities, requiring the learner to have a deeper understanding of the relationships and dependencies between various ontological elements [25, 35, 37]. These advanced question types are particularly valuable for evaluating higher-order thinking and comprehension [12, 25, 37], and whilst MCQs and multi-entity questions may overlap with the previous categories in their use of classes, properties, or instances, their defining characteristic is the integration of multiple facts to form complex questions and the construction of distractors, that introduces an additional cognitive challenge.
2.1 Ontology-based Question Generation
Several studies have explored ontology-based question generation from diverse perspectives by applying these strategies to generate educationally relevant questions. Alsubait et al. [4] proposed a structural method for generating analogy-based MCQs. Their system constructs questions in the form “A is to B as C is to D” by identifying consistent class–subclass paths in ontologies and scoring the structural similarity of candidate pairs. Distractors are selected based on their structural plausibility whilst maintaining a lower analogy strength than that of the correct answer.
@@ 2.1 Ontology-based Question Generation
Generation Strategy RDF Pattern Approaches
@@
Vinu and Kumar [43] introduced a template-driven approach for question generation using ontologies, producing class- and property-based questions based on predefined patterns applied across multiple domain ontologies. Their results demonstrated that the approach was effective in generating a large volume of questions that were semantically comparable to those authored by humans. Cubric and Tosic [12] focused on aligning ontology-based questions with Bloom’s taxonomy [8], by using ontological structures (particularly class, property, and annotation definitions) to generate questions targeting different cognitive levels, including recall, comprehension, and application. In their evaluation, 69% of the generated questions were judged suitable for use in assessments, and 33% were classified as medium to high in cognitive difficulty. The study also reported notable variation in question quality depending on the ontology used, the distractor generation strategy, and the Bloom level targeted; questions that assessed application of knowledge and those based on semantic strategies were rated as highest in quality. More recently, Raboanary et al. [36] proposed a flexible architecture for generating a variety of question types, including yes/no, true/false, subclass, equivalence, and definition questions. Their method relies on axiom-based templates and supports both foundational ontology alignment and natural language realisation. The system was evaluated across multiple ontologies and demonstrated the feasibility of producing a wide range of linguistically fluent and structurally valid questions.
2.2 Ontology Evaluation
Domain ontologies, which represent the knowledge of a specific subject area, form the primary input for the different question generation models discussed above. Many studies rely on hand-crafted ontologies developed specifically for evaluation purposes [2, 5, 13, 24, 43, 42, 40], which frequently focus on the knowledge of a particular course [5, 24], or represent broader knowledge of a general domain [2, 13, 24, 43, 42, 40]. Consequently, such ontologies tend to be small in scope and shallow in structure, capturing only the most general aspects of the domain. This limited depth and scale may constrain the capabilities of question generation models and hinder a comprehensive assessment of their effectiveness. Additionally, the process of building domain-specific ontologies from scratch is time-intensive and requires specialised expertise, posing a challenge to the acceleration of research in this area. Furthermore, whether a new ontology is engineered or an existing one is reused, it is still important to quantitatively assess if it is fit for the task of question generation through the use of appropriate ontology evaluation metrics.
@@ 2.2 Ontology Evaluation
As ontologies are used across diverse application contexts, it is crucial that robust evaluation methods are used to assess them before they are adopted within new systems. Ontology evaluation involves a comprehensive examination of their structure, alignment with domain requirements, and use for specific tasks [18, 19]. However, when reusing an ontology for a specific task, it is imperative to evaluate its suitability and alignment with the task’s requirements. This is particularly important as ontology evaluation depends heavily on the requirements of the task in question.
A requirement is defined as “an expression of desired behaviour” [32], with ontology requirements specifically denoting the desired quality or competency of an ontology within the context of an application [45]. Given the diversity of ontology requirements across applications, a thorough analysis of the target application needs is essential before ontology reuse. Once these requirements are identified, they can be operationalised qualitatively, through expert judgments or user feedback, or quantitatively, employing measurable metrics (i.e. measures). Although various evaluation measures for ontologies have been proposed, selecting appropriate measures should be guided by the application’s requirements rather than relying on an arbitrary set of metrics [16, 27, 31, 45].
@@ 3 Requirement-Oriented Methodology for AQG
Table 2: Basic statistics of the ontologies utilised in the expert-led study.
@@
To determine an appropriate set of ontology-evaluation criteria for AQG, it is necessary to first identify those ontology characteristics that are both relevant to the generation of good quality questions, as well as being measurable. ROMEO (Requirements-Oriented Methodology for Evaluating Ontologies) is a requirement-oriented approach for identifying a set of unique requirements and characteristics for a task domain (in this context, AQG), that can then be aligned to a set of relevant evaluation metrics [45]. It was originally adapted from the Goal-Question-Metrics (GQM) framework [10], designed to derive quality measures by linking them to the goals of a given software application. By determining the requirements of the AQG process through a user-based study, suitable metrics can be identified that quantify the degree to which an ontology can effectively fulfil its intended purpose; i.e. facilitating the generation of high calibre and cognitively challenging questions.
An expert-led study was therefore conducted to determine a number of ontology requirements given a set of generated questions, based on the following stages:
  1. 1.
    A small number of existing domain ontologies were selected for the purpose of generating questions (Table 2); namely, an astronomy-based ontology focused on the solar system, an ontology representing U.S. geographical data, [^2] http://www.cs.utexas.edu/users/ml/geo.html and an ontology capturing intuitive knowledge about African wildlife [20].
  2. 2.
    Several question generation patterns were utilised to generate questions by instantiating templates with ontology elements (Table 1). This resulted in the generation of 3858 questions [^3] For MCQs, only unique question stems were considered; repeated questions with varying answer choices were excluded to avoid redundancy. across all three ontologies (Table 3).
  3. 3.
    An expert familiar with ontology-based AQG was asked to assess the quality of questions, by completing a questionnaire containing a mix of open and closed questions. Two criteria were used to assess the questions: effectiveness and appropriateness (see below). Furthermore, they were asked to identify any ontology-related requirements that may impact the quality of the generation process.
  4. 4.
Table 3: Number of questions generated for each question generation strategy for each ontology in the expert-led study.
Solar System Geography African Wildlife
Class membership 70 713 0
Property-based 120 2044 0
Terminology-based 4 1 25
Annotation-based 10 0 11
MCQs 8 0 0
Multi-entity 17 724 0
Total 351 3471 36
The ontologies listed in Table 2 were chosen as they exhibit different characteristics and thus could be used to examine how their structural differences influence the question generation outcome, such as an emphasis on taxonomic modelling or hierarchical classifications and conceptual relationships. Two of these were populated to provide rich factual and instance-based representations. The questions that were presented to the expert were generated using an AQG approach that utilised the RDF patterns in Table 1 within a fixed textual template. Additional post-processing techniques, such as verbalisation (which are typically carried out to enhance the generation results), were intentionally excluded to maintain consistency in the structure and format of the questions across different ontologies, and to ensure that the expert-led comparison of question quality and effectiveness was not influenced by such techniques. [^4] Example questions in this paper have been paraphrased for clarity. The expert reviewed the question specifications (Table 1) and evaluated the questions generated from each of the three ontologies, with the aim of identifying specific ontological features that contributed to variations in the questions. The two evaluation criteria employed by Leo et al. [25] were used when assessing each ontology and its corresponding questions:
@@
  • The question quality (appropriateness) is assessed based on how pedagogically appropriate the questions are, to assess learners at different knowledge levels by ensuring questions range from basic to a more advanced understanding of the domain.
Table 4: Symbols used by the metrics following the nomenclature used by Tartir et.al. [38, 39].
@@
$\mathcal{H}^{\mathcal{C}}$ a concept hierarchy, which is a directed, transitive relation, and relates to the taxonomy. $\mathcal{H}^{\mathcal{C}}(c_{1},c_{2})$ states that $c_{1}$ is a sub-concept of $c_{2}$ (i.e. $c_{1}\subseteq c_{2}$).
propprop: $\mathcal{P}\to\mathcal{C}\times\mathcal{C}$: A function mapping properties that link pairs of concepts. $prop(p)=(c_{1},c_{2})$ can also be expressed as $p(c_{1},c_{2})$.
$\mathcal{A}$ a set of annotation properties of the type rdfs:comment.
@@
$inst^{\mathcal{C}}$: $\mathcal{C}\to 2^{\mathcal{I}}$ (concept instantiation) maps concepts to their instances. $inst^{\mathcal{C}}(c)=i$ may also be written as $c(i)$.
$inst^{\mathcal{P}}$: $\mathcal{P}\to 2^{\mathcal{I}\times Ob}$. For each property $p\in\mathcal{P}$, $inst^{\mathcal{P}}$ is the set of pairs $(i,o)$ such that the assertion $p(i,o)$ holds, with $i\in\mathcal{I}$ and $o\in Ob=\mathcal{I}\cup L$. This may also be written as $p(i,o)$.
As this evaluation task involves an extensive qualitative analysis of a large pool of questions generated from three diverse ontologies, a single expert was recruited by this study to assess the feasibility of the proposed approach, as well as to understand critical aspects such as the time and effort needed by experts for such evaluation. The recruitment of a singe expert was sufficient in this case as the objective was to identify structural and semantic features of ontologies through pedagogically relevant questions, rather than to establish consensus or benchmark performance.
4 Response Analysis and Requirement Derivation
This section presents the outcomes of the expert-led study, by describing the observations made by the expert, the requirements that the expert derived from these observations, and proposed metrics that were subsequently defined to measure these requirements. The proposed evaluation metrics build upon well-established measures [16, 28, 38, 39] that have been adapted to align with the specific requirements of the AQG task. The model used for defining metrics formally specifies the ontology schema and the knowledge base structures. An ontology schema represents the class and property definitions (i.e. T-box) and is denoted by the 5-tuple $\mathcal{O}=\{\mathcal{C},\mathcal{P},\mathcal{H}^{\mathcal{C}},prop,\mathcal{A}\}$, whereas a knowledge base (i.e. T-box $\cup$ A-Box) represents the instantiated ontological elements (individuals) and is denoted by the 5-tuple $\mathcal{KB}=\{\mathcal{O},\mathcal{I},L,inst^{\mathcal{C}},inst^{\mathcal{P}}\}$. The meaning of the symbols used in these definitions are given in Table 4. Whilst this model is similar to others proposed for various ontology scenarios [28], it has been adapted to be more pertinent to the metrics that focus on AQG tasks.